A plant requires approximately 4178 kcal for the production of 1.00 kg of starch from carbon dioxide and water.


Calculate the energy in kilocalories required by a plant for the production of 7.69 g of starch.

Energy=.............kcal


Please with steps



1
Expert's answer
2021-03-02T01:02:36-0500

1.00 Kg of starch = 1000 g of starch.

Given, to produce 1000 g of starch the required energy is 4178 KCal.

Therefore, to produce 1 g of starch the required energy = 4178/1000 KCal = 4.178 KCal.

So, for the production of 7.69 g of starch the required energy = 4.178 × 7.69 KCal = 32.13 KCal.